OpenStudy (anonymous):

1 and two-fifths times 4 and two-thirds equals

OpenStudy (anonymous):

\[1\frac{2}{5}*4\frac{2}{3}=?\]

OpenStudy (anonymous):

is this it?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

basically set the whole thing as just two fraction with nothing next to it.

OpenStudy (anonymous):

1 and two-fifth = 1 2/5 4 and two-third = 4 2/3 Converting mixed fractions to improper fractions, 1 2/5 = (1*5+2) / 5 = 7/5 4 2/3 = (4*3+2) / 3 = 14/3 Multiplying them, we get \[\huge 7/5 * 14/3 = (7 * 14) / (5 * 3) = 98/15 \] Therefore the answer is \[\huge 98/15\]

OpenStudy (anonymous):

Next time try not to give the answer

OpenStudy (anonymous):

make the asker do a little bit of the work

OpenStudy (anonymous):

why so I am here to help not sleep

OpenStudy (anonymous):

I don't think you help people by giving answers....